OpenStudy (anonymous):

Housework study The New York State College of Human Ecology at Cornell University conducted a study of housework. The study found that the average stay-at-home mom puts 8 hours a day into housework. What about the husband? He does only 96 minutes a day! Here is what the husband does in his 96 minutes of chores: Cleaning 36 minutes Kitchen work 12 minutes Caring for the family 24 minutes Shopping and paperwork 24 minutes What reduced fraction represents the husband's kitchen work?

OpenStudy (kropot72):

Total time per day spent by the husband on housework = 96 minutes Time per day spent by husband on kitchen work = 12 minutes Fraction of husband's housework time spent on kitchen work = 12/96 = 1/8

OpenStudy (anonymous):

tyvm!

OpenStudy (kropot72):

You're welcome :)